OpenStudy (cute124etuc):

F (x) = 9x^3 + 2x^2 - 5x+4 and g(x) = 5x^3 - 7x + 4 what is F (x) - g(x)

OpenStudy (cute124etuc):

@AloneS can you help

OpenStudy (cute124etuc):

@jhonyy9 can you help

jhonyy9 (jhonyy9):

just you need assuming or subtracting the same terms - ok. ?

OpenStudy (cute124etuc):

i dont understand can you walk me through it step by step

jhonyy9 (jhonyy9):

like the first term 9x^3 -5x^3 = ?

OpenStudy (cute124etuc):

4x^3 is that correct

jhonyy9 (jhonyy9):

yes is right sure

jhonyy9 (jhonyy9):

so than can you continue it ?

OpenStudy (cute124etuc):

I will try solving it and type it my answer can you check it for me when i am done

jhonyy9 (jhonyy9):

yes sure but rewrite it here please now what about 2x^2 ?

jhonyy9 (jhonyy9):

so bc. inside g(x) havnt this term of x^2 so this will remain it in this style

OpenStudy (cute124etuc):

5x^3 - 7x+4 = ?

jhonyy9 (jhonyy9):

-5x-(-7x)=?

OpenStudy (cute124etuc):

2

OpenStudy (cute124etuc):

but where did the 4 at the end go

jhonyy9 (jhonyy9):

there above is got 4x^3 not 5

jhonyy9 (jhonyy9):

you have said me step by step so than please

jhonyy9 (jhonyy9):

so are you here ?

OpenStudy (cute124etuc):

@jhonyy9 i am so sorry my wifi keeps crashing

jhonyy9 (jhonyy9):

-5x-(-7x)=?

OpenStudy (cute124etuc):

2x

OpenStudy (cute124etuc):

is that correct ?

OpenStudy (cute124etuc):

@jhonyy9 are you here?

OpenStudy (cute124etuc):

is the final answer for the whole thing 2^3 + 4 @jhonyy9

jhonyy9 (jhonyy9):

the first term will 4x^3 +2x^2 -5x-(-7x) ...

jhonyy9 (jhonyy9):

so you need to get -5x-(-7x) = ?

OpenStudy (cute124etuc):

one second i think i am starting to get it can i try to solve just give me a sec

OpenStudy (cute124etuc):

4x^3 + -5xx + 5x

OpenStudy (cute124etuc):

@jhonyy9 is it correct I meant 4x^3 + -5x + 5x

jhonyy9 (jhonyy9):

no the first two terms are 4x^3 +2x^2 -5x-(-7x) ... so -5x-(-7x)= ?

OpenStudy (cute124etuc):

2x

OpenStudy (cute124etuc):

is that correct ?

jhonyy9 (jhonyy9):

yes -5x-(-7x)=-5x+7x = 2x

OpenStudy (cute124etuc):

then ?

jhonyy9 (jhonyy9):

+4-(+4) = ?

OpenStudy (cute124etuc):

0

jhonyy9 (jhonyy9):

yes exactly

OpenStudy (cute124etuc):

so what is the answer 0 ?

OpenStudy (cute124etuc):

@jhonyy9

OpenStudy (cute124etuc):

is 0 the correct answer ?

jhonyy9 (jhonyy9):

yes at the last one so than your answer will be how ?

jhonyy9 (jhonyy9):

4x^3 +2x^2 -5x-(-7x) ... 4x^3 +2x^2 +2x + ?

OpenStudy (cute124etuc):

so what is the final answer liek f (x) - g (x) = ?????

jhonyy9 (jhonyy9):

4x^3 +2x^2 +2x+0 so than will be 4x^3 +2x^2 +2x

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (cute124etuc):

yes thank you

jhonyy9 (jhonyy9):

was my pleasure hope you understand it now bye bye